OpenStudy (anonymous):

A school ordered $600 worth of lightbulbs. Some of the lightbulbs cost $1 each and the other cost $2 each. if twice as many $1 bulbs as $2 bulbs were ordered, how many lightbulbs were ordered altogether?

OpenStudy (anonymous):

Please show how you got your answer!

OpenStudy (anonymous):

laptop is slow put x = number of $2 light bulbs then since there are twice as may $1 light bulbs as $2 light bulbs you know that the number of $1 light bulbs is \[2x\] and then you get \[2x\times $1 +x\times $2 = $600\]

OpenStudy (anonymous):

\[4x=600\] \[x=600\div 4\] \[x=150\] so 150 light bulbs that cost $2 and 300 light bulbs that cost $1